OpenStudy (anonymous):

Help, will medal and fan if given correct answer!! Tracy need to find the length and width of the equipment storage room where she works. The room is 1 meter longer than it is wide. Its perimeter is 14 meters. Find the width in meters.

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

2(w+w+1) = 14 => 4w +2 = 14 or w = 3

OpenStudy (anonymous):

as room is ractangle perimeter = 2*(l+b) given in question is l = b+1 so Perimeter = 2((b+1) +b) P = 2(2b +1) 14 = 2(2b +1) 7 = 2b +1 6 = 2b b = 3 l = b+1 l = 3+1 l = 4